# ProbRadarM3F: mmWave Radar-based Human Skeletal Pose Estimation with Probability Map Guided Multi-Format Feature Fusion

### v2

@@ III-C Multi-format Feature Fusion and Estimation Head
where $i$ indicates the layers after 3D convolution. $F_{f}$ represents the fused feature of each layer, and $F_{1}$, $F_{2}$ represent features from two branches.
@@
The predicted keypoint heatmaps are obtained through the attention decoding module from the multi-layer features. Knowing the location of certain keypoints can help in predicting the location of other keypoints. Therefore, similar to [16], the pose refinement module of the Graph Convolutional Network (GCN) is used to refine the keypoint heatmap [35]. A 3-layer GCN is used to perform feature propagation and inference, refining the keypoint predictions using the mutual position information of the physical connections between keypoints. The generated new heatmap is used as a common reference to generate the final keypoint locations. There is a mismatch between the coordinate systems for radar signals and keypoint heatmaps, making it impossible to directly use keypoint coordinates to locate radar features.
To facilitate the end-to-end training of the model, a strategy of imposing a pixel-wise binary cross-entropy loss on both the initial keypoint heatmaps and the GCN-refined keypoint heatmaps is employed. The objective function is formulated as follows:
@@
where ${H}$ and ${\hat{H}}$ represent the initial keypoint heatmaps and refined keypoint heatmaps, respectively, and ${G}$ represents the generated groundtruth keypoint heatmaps based on Gaussian distribution. To illustrate, the binary cross-entropy loss $L_{\mathrm{bce}}(H,G)$ is defined by the equation:
@@
where $c,i,j$ is derived from the channel, width and height of the joint prediction heatmaps $H$ and groudtruth heatmaps $G$, which follow Gaussian distribution.
IV Experiments and Results
@@
IV-A Dataset and Evaluation
The HuPR dataset was chosen for our experiments [16]. This dataset acquires data from two identical radars. One radar sensor is rotated 90° in the antenna plane with respect to the other, with one radar focusing on the horizontal plane and the other on the vertical plane. Unlike some radar datasets that provide processed data in formats such as point clouds [36, 37, 38], heatmaps [39], etc., which lose the purity of radar data, HuPR provides raw radar ADC signals. This allows more possibilities for our data processing and feature extraction methods.
@@ IV-A Dataset and Evaluation
In alignment with established norms in the field, our evaluation framework for 2D keypoints employs average precision (AP) metrics calculated across various levels of object keypoint similarity (OKS) [41]. The OKS is calculated using the following formula:
@@
This allows the evaluation of the model’s accuracy in detecting 2D keypoints across 14 critical human body positions, including the head, neck, shoulders, elbows, wrists, hips, knees, and ankles. To offer a nuanced understanding of the model’s performance, three distinct AP metrics are used: $AP^{50}$, $AP^{75}$, and $AP$. These metrics represent varying degrees of OKS stringency, with $AP^{50}$ and $AP^{75}$ indicating more lenient and stringent OKS constraints, respectively. The metric denoted as $AP$ calculates the mean average precision across a range of 10 OKS thresholds, specifically at 0.5, 0.55, and incrementally up to 0.95, providing a comprehensive overview of the model’s performance across a broad spectrum of precision requirements.
IV-B Implementation Details
We take 193 sequences from HuPR for training, 21 for validation, and 21 for test. The sequence are chosen as HuPR network, which is considered as baseline, for fair comparison of results.
We implement ProbRadarM3F using Python and Pytorch. The training of our network is processed on a single NVIDIA V100 GPU. The network is trained using the step decay learning rate strategy, with an initial learning rate of 0.00005 and a reduction of 0.999 times per 2000 iterations. The Adam is employed as optimizer, and we set the batch size to 24. Some other parameters in the experiment are set as follows. In CFAR processing, the settings of the guard unit and the reference unit are set to 16 and 8. The depth of positional encoding for the probability map is set to 32. The number of input frames is 8 for joint processing of multi-frame data. The convolutional of network contains basic blocks from ResNet [42]. The size of convolution kernel for 3D convolutional processing is $3\times 3\times 3$ and ReLU activation function is used. These parameters are selected based on optimal performance in experiments.
IV-C Results and Analysis
[…]
Fig. 4 illustrates the performance of our model under different types of actions in comparison. GT represents the ground truth keypoints generated from HRNet [40] that precisely follow the actions in RGB frames, while HuPR and Ours are the predictions of the baseline and our ProbRadarM3F for the same frame, respectively. As shown, the predicted keypoints mostly align with the ground truth. To present realistic results, some frames where the prediction is not accurate enough are displayed in Fig. 5. It can be seen that our model loses accuracy when the target moves quickly. Although it predicts the torso and head locations well, there is still room for improvement in the wrist and foot joints. This observation highlights a critical area for future improvement, particularly in enhancing the model’s sensitivity to high-motion extremities.
Therefore, analyzing both the overall and individual results, our work effectively improves the precision of mmWave radar-based human skeletal pose estimation. This demonstrates that the probability map-guided positional encoding method effectively mines information from radar signals that might have been previously overlooked.
TABLE I: Comparison of keypoint Accurate Precison
@@ IV-C Results and Analysis
TABLE II: Comparison of State-of-the-art Approaches and Our Proposed Model
@@
IV-D Ablation Study
@@
  • $\bullet$
[…]
  • $\bullet$
    In addition to isolating the effect of the ProbPE branch as a whole, the influence of probability maps in the ProbPE branch was specifically assessed. The complete model conducts positional encoding guided by probability maps, which are pivotal in generating refined positional features. To quantify the contribution of these probability maps, an ablation experiment was conducted by excluding the probability map generation step and directly applying positional encoding to the data after DOA. As shown in Table III, precision improved compared to using the FFT branch independently. However, the absence of probability maps resulted in a noticeable decrease in performance, leading to a reduction of $2.3\%$ in $AP$ when compared to the complete model. The results demonstrate the importance of probability maps. Without the guidance of probability maps, positional encoding becomes less effective. Consequently, the probability maps are not merely auxiliary but integral to ensuring that the positional encoding maximally captures and utilizes the spatial context inherent in the radar signals. The probability maps enhance the model’s ability to accurately determine the probability of target presence at specific azimuth and elevation coordinates, which is crucial for accurately encoding positional information.
TABLE III: Ablation Study
@@
V Conclusion
In our paper, we have explored the use of mmWave radar for estimating human skeletal poses. This is an important component of smart sensor systems used in medical care IoT applications, particularly in addressing privacy concerns. We introduced ProbRadarM3F, a new model that uses millimeter-wave radar for human skeletal pose estimation. This model significantly improves the extraction and use of hidden information from radar signals. We also introduced the ProbPE branch, which generates probability maps based on estimating the likelihood of a specific position. These probability maps are then used to effectively extract position features using a positional encoding method. Additionally, we combined positional features with features obtained from the FFT branch to enhance the human body keypoint features produced by the model. Our experiments on the HuPR dataset demonstrated that ProbRadarM3F outperforms existing methods, indicating the effectiveness of our approach. Our results highlight that there is valuable information, such as positional data, present in the radar heatmap or even in the radar signal itself, which should not be considered redundant.

### v3

@@ 3.3 Multi-format Feature Fusion and Estimation Head
where $i$ indicates the layers after 3D convolution. $F_{f}$ represents the fused feature of each layer, and $F_{1}$, $F_{2}$ represent features from the two branches, respectively.
@@
The predicted keypoint heatmaps are obtained through the attention decoding module from the multi-layer features. Knowing the location of certain keypoints can help in predicting the location of other keypoints. Therefore, same as [18], the pose refinement module of the Graph Convolutional Network (GCN) is used to refine the keypoint heatmap [40]. A 3-layer GCN is used to perform feature propagation and inference, refining the keypoint predictions using the mutual position information of the physical connections between keypoints. The generated new heatmap is used as a common reference to generate the final keypoint locations. There is a mismatch between the coordinate systems for radar signals and keypoint heatmaps, making it impossible to directly use keypoint coordinates to locate radar features.
To facilitate the end-to-end training, a strategy of imposing a pixel-wise binary cross-entropy loss on both the initial keypoint heatmaps and the GCN-refined keypoint heatmaps is employed. The objective function is formulated by
@@
where ${H}$ and ${\hat{H}}$ represent the initial keypoint heatmaps and refined keypoint heatmaps, respectively, and ${G}$ represents the generated ground truth keypoint heatmaps based on Gaussian distribution. To illustrate, the binary cross-entropy loss $L_{\mathrm{bce}}(H,G)$ is defined by
@@
where $c,i,j$ are derived from the channel, width and height of the joint prediction heatmaps $H$ and groud truth heatmaps $G$, which follow a Gaussian distribution.
4 EXPERIMENTS AND RESULTS
@@
This section introduces the implementation and detailed performance of our proposed method. Ablation experiments are also constructed to compare the prediction performance of incomplete models.
4.1 Dataset and Evaluation
The HuPR dataset was chosen for our experiments [18]. This dataset acquires data from two identical radars. One radar sensor is rotated $90^{\circ}$ in the antenna plane with respect to the other, where one radar focuses on the horizontal plane and the other on the vertical plane. Unlike some radar datasets that provide processed data, e.g., point clouds [41, 42, 43], heatmaps [44], etc., which lose the purity of radar data, HuPR provides raw radar ADC signals. Table 1 demonstrates a comparison of the data formats provided by mainstream accessible mmWave radar-based human posture datasets. HuPR dataset the only dataset that provides radar signals that have not been manually processed. This allows for more possibilities for our data processing and feature extraction methods.
Table 1: Data Formats for MmWave Radar Human Pose Datasets
$Dataset$ $Radar$ $RawSignal$ $PointCloud$ $Heatmap$
Mars[41] TI IWR1443 $\circ$ $\bullet$ $\circ$
mRI[45] TI IWR1443 $\circ$ $\bullet$ $\circ$
mmPose[27] TI AWR1642 $\circ$ $\bullet$ $\circ$
miliPoint[42] TI IWR1843 $\circ$ $\bullet$ $\circ$
HuPR[18] TI IWR1843 $\bullet$ $\circ$ $\circ$
mmPose-FK[46] TI IWR6843 $\circ$ $\bullet$ $\circ$
mmBody[47] Arbe Robotics Phoeniex $\circ$ $\bullet$ $\circ$
HIBER[30] TI MMWCAS-RF-EVM $\circ$ $\circ$ $\bullet$
@@ 4.1 Dataset and Evaluation
In alignment with established norms in the field, our evaluation framework for 2D keypoints employs average precision (AP) metrics calculated across various levels of object keypoint similarity (OKS) [49]. The OKS is calculated by
@@
This allows the evaluation of model accuracy in detecting 2D keypoints across 14 critical human body positions, including head, neck, shoulders, elbows, wrists, hips, knees, and ankles. To offer a nuanced understanding of model performance, three distinct AP metrics are used: $AP^{50}$, $AP^{75}$, and $AP$. These metrics represent varying degrees of OKS stringency, with $AP^{50}$ and $AP^{75}$ indicating more lenient and stringent OKS constraints, respectively. The metric $AP$ calculates the mean average precision across a range of 10 OKS thresholds, specifically at 0.5, 0.55, and incrementally up to 0.95, providing a comprehensive overview of the model’s performance across a broad spectrum of precision requirements.
4.2 Implementation Details
We take 193 sequences from the HuPR dataset for training, 21 for validation, and 21 for testing. The sequences are chosen as HuPR network, which is considered as baseline, for fair comparison of results.
ProbRadarM3F is implemented using Python and Pytorch. The training of our network is processed on a single NVIDIA V100 GPU. The network is trained using the step decay learning rate strategy, with initial learning rate $0.00005$ and reduction $0.999$ times per $2000$ iterations. The Adam is employed as optimizer, and the batch size is set to $32$. Some other parameters in the experiment are set as follows. In CFAR processing, the settings of the guard unit and the reference unit are set to 5 and 16. The depth of positional encoding for the probability map is set to 32. Table 2 shows how the model performs for some of the keypoints and the overall prediction for different positional encoding layers. The number of input frames is 8 for joint processing of multi-frame data. The convolution network contains basic blocks from ResNet [50]. The size of convolution kernel for 3D convolutional processing is $3\times 3\times 3$ and the ReLU activation function is used. These parameters are selected based on optimal performance in experiments.
Table 2: Comparison of Different Positional Encoding Layers
$Layers$ AP
$Head$ $Elbow$ $Wrist$ $Total$
8 80.9 48.8 27.0 69.2
16 82.9 50.9 26.9 68.3
32 81.1 52.2 28.2 69.9
64 80.0 45.8 23.5 0.65
Table 3: Comparison of keypoint Accurate Precison
@@ 4.2 Implementation Details
Table 4: Comparison of State-of-the-art Approaches and Our Proposed Model
@@
4.3 Results and Analysis
Experiments on the HuPR dataset were performed to evaluate the effectiveness of the proposed strategies.
[…]
Fig. 4 illustrates the performance of our model under different types of actions. GT represents the ground truth keypoints generated from HRNet [48] that precisely follow the actions in RGB frames, while HuPR and Ours are the predictions of the baseline and our ProbRadarM3F for the same frame, respectively. As shown, the predicted keypoints mostly align with the ground truth. To present realistic results, some frames where the prediction is not accurate enough are displayed in Fig. 5. It can be seen that our model loses accuracy when the target moves quickly. Although it predicts the torso and head locations well, there is still room for improvement in the wrist and foot joints. This observation highlights a critical area for future improvement, particularly in enhancing the model’s sensitivity to high-motion extremities.
Therefore, by analyzing both the overall and individual results, our work effectively improves the precision of mmWave radar-based human skeletal pose estimation. It demonstrates that the probability map-guided positional encoding method effectively mines information from radar signals that might have been previously overlooked.
4.4 Ablation Study
@@
Table 5: Ablation Study
@@
  • $\bullet$
[…]
  • $\bullet$
    In addition to isolating the effect of the ProbPE branch as a whole, the influence of probability maps in the ProbPE branch was specifically assessed. The complete model conducts positional encoding guided by probability maps, which are pivotal in generating refined positional features. To quantify the contribution of these probability maps, an ablation experiment was conducted by excluding the probability map generation step and directly applying positional encoding to the data after DOA. As shown in Table 5, precision improved compared to using the FFT branch independently. However, the absence of probability maps resulted in a noticeable decrease in performance, leading to a reduction of $2.3\%$ in $AP$ when compared to the complete model. The results demonstrate the importance of probability maps. Without the guidance of probability maps, positional encoding becomes less effective. Consequently, the probability maps are not merely auxiliary but integral to ensuring that the positional encoding maximally captures and utilizes the spatial context inherent in the radar signals. The probability maps enhance the model’s ability to accurately determine the probability of target presence at specific azimuth and elevation coordinates, which is crucial for accurately encoding positional information.
5 CONCLUSION
In this paper, mmWave radar was applied to estimate human skeletal poses. This is an important component of smart sensor systems used in medical care IoT applications, particularly in addressing privacy concerns. ProbRadarM3F is designed to apply millimeter-wave radar for human skeletal pose estimation. This model significantly improves the extraction and use of hidden information from radar signals, by introducing the ProbPE branch, which generates probability maps based on estimating the likelihood of a specific position. These probability maps are then used to effectively extract position features using a positional encoding method. Additionally, we combined positional features with features obtained from the FFT branch to enhance the human body keypoint features produced by the model. Our experiments on the HuPR dataset demonstrated that ProbRadarM3F outperforms existing methods, indicating the effectiveness of our approach. Our results highlight that there is valuable information, such as positional data, present in the radar heatmap or even in the radar signal itself, which should not be considered redundant in human pose estimation tasks.